Jimmy invest Rm 15,000 in an account for six years. The investment account pays 8%
compounded semi-annually for the first two years and 9% compounded monthly for the
rest of the period.
a. What is the maturity value of this investment?
b. Find the interest earned from this investment.
2. Given two interest rate (a) 4.8% compounded monthly and (b) 5% compounded every
three months, which of the two interest rates provides a better return?

Answer by Boreal(15235)

P=Po(1+(r/t))^nt
15000(1+.08/2)^4, divide the interest rate by 2 for the compounding, and multiply the number of compoundings by 2 because semiannual.
=$17547.88
Now use that for 9% compounded monthly
P=17547.88(1+(.09/12))^48, the 48 being 12 monthly compoundings for 4 years: $25,228.13, rounding at the end. MATURITY VALUE
The interest earned:
first two years: $2547.88
Total: $10,228.13-$2547.88 earned in two years=
7680.25 earned in last four years
TOTAL INTEREST IS $10228.13
================================
Do it for one year
(1+(.048/12))^12=1.04907
(1+.05/4)^4=1.05094
The second one, at 5% every 3 months.